OpenStudy (anonymous):

I need help with this logarithm equation on the steps to solve. 400/1+e^-x=350

OpenStudy (anonymous):

400 /(1+e^x) = 500?

OpenStudy (anonymous):

but not 500 , 350

OpenStudy (anonymous):

oh yes - sorry

OpenStudy (anonymous):

I got -0.13353192

OpenStudy (anonymous):

Multiply through by the denominator to get 400=350+350e^x Then isolate the variable: 50/350=e^x Apply the definition of logarithm: x=ln(1/7)

OpenStudy (anonymous):

I have the answer, thas not it . @ebnor

OpenStudy (anonymous):

\[400/(1+e ^{-x}) = 350\]\[400/1=350e ^{-x}\]\[400/350 = e ^{-x}\]\[\ln 400/350= lne ^{-x}\]\[-0.13353192 =x\] Where did I go wrong?

jimthompson5910 (jim_thompson5910):

\[\large \frac{400}{1+e^{-x}}=350\] \[\large 400=350(1+e^{-x})\] \[\large 400=350+350e^{-x}\] \[\large 400-350=350e^{-x}\] \[\large 50=350e^{-x}\] \[\large \frac{50}{350}=e^{-x}\] \[\large \frac{1}{7}=e^{-x}\] \[\large \frac{1}{7}=\frac{1}{e^{x}}\] \[\large e^{x}=7\] \[\large x=\ln(7)\]

OpenStudy (anonymous):

The answer is suppose to be 1.95491.

OpenStudy (anonymous):

@jim explained it logically , thanks

jimthompson5910 (jim_thompson5910):

The exact answer is \[\large x=\ln(7)\] the approximate answer is \[\large x\approx 1.94591014\]

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

for the ones with negative answers, they lost a negative somewhere

OpenStudy (anonymous):

Yeah, I redid it with a few changes and got that too. I didn't carry the "1" over in my first attempt, that's where I went wrong. Can someone help me with my question?

jimthompson5910 (jim_thompson5910):

post to the left (if the question is completely different) or ask here (if it's the same or very similar)

OpenStudy (anonymous):

woops didnt see the negative x. so i redid it above with a -x

jimthompson5910 (jim_thompson5910):

personally, it's better to start a new thread since large threads tend to lag

OpenStudy (anonymous):

I've already started one